Archaeologists find pottery fragments and shell beads at an ancient site. There are 87 objects in all. There are 65 more pottery fragments than shell beads. What system of linear equations will represent this scenario?
Answer by ikleyn(52788)

.
Archaeologists find pottery fragments and shell beads at an ancient site. There are 87 objects in all.
There are 65 more pottery fragments than shell beads. What system of linear equations will represent this scenario?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

Let x = the number of pottery fragments 
and y = the number of shell beads.

Then

x + y = 87,
x = y + 65.

or (which is the same)

x + y = 87,
x - y = 65.